How do you solve functions? for exampl:f(x)=3x+2
and f(x)=x/x-1

Answer by prince_abubu(198)

Actually, you DON'T solve functions. In beginning Algebra, they give you the function definition and they ask you, for example, to find the value of f(7). What this means is for you to plug in 7 in the expression to the right of the equals sign. So f(7) = 3(7) + 2 = 21 + 2 = 23. So you can then write it straight-up that f(7) = 23.

They may also ask you to graph f(x) = 3x + 2. That's the same thing as graphing y = 3x + 2. The f(x) is just a fancier way of saying y.

The whole idea behind functions is to get you to think "I put in something, I get something else out" or "a specific action leads to a specific consequence". It happens a lot in real life.